OpenStudy (anonymous):

Write this on vertex form y=x^2+14x+49-49

OpenStudy (anonymous):

y = (x² - 14x + 7²) + 30 - 7² y = (x - 7)² + 30 - 49 y = (x - 7)² - 19 i think its right..

OpenStudy (anonymous):

if i think i know what you said..